Phobos is the closer of Mars' two small moons, has
Answer #1

R =9400 km , M =6.4*10^23 kg , G =6.67*10^-11 N .m^2/kg^2

From Keplers third law

T^2 = (4pi^2/GM)R^3

T^2 = (4*3.14*3.14*(9400*1000)^3)/(6.67*10^-11*6.4*10^23)

T =27701.2 s

T = 7.695 hours